### v1

@@ II $\Xi(1530)$ production in the $Kp$ scattering process
Figure 2: Diagrams contributing to the process of $K^{-}p\rightarrow K^{0}\Xi(1530)^{0}$, in which (a) and (b) correspond to the $s$- and $u$-channel contributions, respectively.
In the same way, one can investigate the $\Xi(1530)^{0}$ production in the $K^{-}p$ scattering process, and the relevant diagrams are shown in Fig. 2. It should be noted that in the $u$-channel, the exchanged hyperons should only be $\Sigma$ and its resonances, while $\Lambda$ hyperon and its resonances are forbidden due to charge conservation. Thus, from the isospin symmetry, the contributions from $\Sigma$ hyperon and its resonances are the same in both $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$ processes, while those from $\Lambda$ hyperon and its resonances are different in both processes.
III NUMERICAL RESULTS AND DISCUSSIONS
To describe the experimental data of the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$ Berge:1966zz ; Dauber:1969hg ; Briefel:1977bp ; Flaminio:1979iz , we have considered nine $\Lambda/\Sigma$ hyperon and their resonances as intermediate states and the involved states are presented in Table 1. For the resonances with high-spin, i.e., $\Lambda(2110)$ and $\Sigma(2250)$, only $s$-channels are considered as in Ref. Sharov:2011xq . Therefore, the amplitude for $K^{-}p\to K^{+}\Xi(1530)^{-}$ reads,
@@ III NUMERICAL RESULTS AND DISCUSSIONS
With the above amplitudes, one can obtain the differential cross sections for the $K^{-}p\to K\Xi(1530)$, which is,
@@
where $s$ refers to the square of the center-of-mass energy, and $\theta$ is the scattering angle, which is the angle of outgoing $\Xi(1530)$ and the incident kaon beam direction in the center-of-mass frame. The $\vec{q_{f}}$ and $\vec{q_{i}}$ stand for three-momentum of the final $\Xi(1530)$ and the initial kaon beam in the center-of-mass system, respectively, and their values are presented in Eqs. (31).
@@
where $Y$ refers to the $\Lambda/\Sigma$ hyperon and their resonances. In the present estimations, the values of $g_{Y}$ will be determined by fitting the cross sections for $K^{-}p\to K^{+}\Xi(1520)^{-}$ and $K^{-}p\to K^{0}\Xi(1520)^{0}$
Figure 3: (Color online) The cross sections for the $K^{-}p\rightarrow K^{+}\Xi(1530)^{-}$ (diagram (a)) and $K^{-}p\rightarrow K^{0}\Xi(1530)^{0}$ (diagram (b)) processes depending on the $\sqrt{s}$. The black points with error bars correspond to the experimental data for $K^{-}p\to K\Xi(1530)$ from Ref. Berge:1966zz ; Dauber:1969hg ; Briefel:1977bp ; Flaminio:1979iz , while the blue solid curves are obtained with the fitted parameters in Table 2.
@@
Figure 5: (Color online) The individual contributions from different intermediate states for the cross sections for $K^{-}p\rightarrow K^{0}\Xi(1530)^{0}$ depending on the $\sqrt{s}$ in Model A.
@@ III.1 Cross sections for $K^{-}p\to K\Xi(1530)$
In the present fitting process, nine $\Lambda/\Sigma$ hyperon and their resonances are taken into consideration. In addition, two kinds of cutoff constants, $\Lambda_{a}$ and $\Lambda_{b}$, are involved by the form factors in the amplitudes, with for low-spin intermediate states ($J^{P}\leq 3/2$) and $\Lambda_{b}$ for high-spin intermediate states ($J\geq 5/2$), respectively. Here, we take $\Lambda_{a}=1.0$ GeV and $\Lambda_{b}=0.5$ GeV as in Ref. Sharov:2011xq , while the rest parameters $g_{Y}$ can be determined by fitting the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$. In addition, it should be noted that the experimental data for the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ in the range of $\sqrt{s}=[2.087,2.168]$ GeV are inconsistent with each other as shown in Fig. 3. In particular, the cross sections at $2.087$ GeV, $2.111$ GeV, $2.123$ GeV, $2.126$ GeV, $2.151$ GeV, and $2.168$ GeV are around 80 $\mathrm{\mathrm{\mu}b}$ with lage uncertainties, while the cross section at $2.108$ GeV is measured to be $(41\pm 10)\ \mathrm{\mu b}$, which is evidently smaller than the nearby data. Due to the inconsistency mentioned above, in the following, we perform the fitting with the experimental data having the same weight (model A) and different weight (model B) .
Table 2: The fitted parameters of the process $K^{-}p\rightarrow K\Xi(1530)$, where the weight of each experimental data remains the same.
Coupling constant Value Coupling constant Value
$g_{\Lambda(1116)}$ $1.373^{+0.196}_{-0.196}$ $g_{\Sigma(1193)}$ $-3.178^{+0.068}_{-0.068}$
$g_{\Lambda(1405)}$ $-0.046^{+0.080}_{-0.080}$ $g_{\Sigma(1670)}$ $-1.545^{+0.195}_{-0.195}$
$g_{\Lambda(1890)}$ $0.496^{+0.267}_{-0.267}$ $g_{\Sigma(2080)}$ $-0.864^{+0.198}_{-0.198}$
$g_{\Lambda(2110)}$ $0.453^{+0.022}_{-0.022}$ $g_{\Sigma(2250)}$ $0.209^{+0.033}_{-0.033}$
$g_{\Lambda(2325)}$ $-0.182^{+0.017}_{-0.017}$
@@
Figure 6: (Color online) The same as Fig. 3 but for the weighted fit. The points with error bars correspond to the experimental data for $K^{-}p\to K\Xi(1530)$ from Ref. Berge:1966zz ; Dauber:1969hg ; Briefel:1977bp ; Flaminio:1979iz , while the blue solid curves are obtained with the fitted parameters in Table 3. The green data are excluded in the present fit, while the weight of the six red data in the range of $\sqrt{s}=[2.087,2.168]$ GeV are set to 8.
@@ III.1.1 Model A
In this scenario, the parameters $g_{Y}$ determined by fitting the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$ are collected in Table 2. With the central values of the determined parameters, one has $\chi^{2}/d.o.f=63.38/50=1.27$, which is acceptable for the present fit. In Fig. 3, we present the fitted cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ (digram (a)) and $K^{-}p\to K^{0}\Xi(1530)^{0}$ (diagram (b)), where the experimental data from Refs. Berge:1966zz ; Dauber:1969hg ; Briefel:1977bp ; Flaminio:1979iz are also presented for comparison. From the figure one can find that the cross sections for $K^{-}p\to K^{0}\Xi(1530)^{0}$ (diagram (b)) can be well reproduced in the whole considered range. Meanwhile, the cross sections for $K^{-}p\to K^{0}\Xi(1530)^{0}$ can also be reproduced except for the data in the range of $\sqrt{s}=[2.087,2.168]$ GeV due to the inconsistency of the experimental data. Particularly, the cross section for $K^{-}p\to K^{+}\Xi(1530)^{-}$ is fitted to be 50.41 $\mu$b at $\sqrt{s}=2.11$ GeV, which is consistent with the experimental measurement within uncertainty. However, the fitted cross sections in the range of $\sqrt{s}=[2.087,2.168]$ GeV are in whole smaller than other data in this range. Furthermore, it is worth noting that there are two clear peak structures around $\sqrt{s}=2.1$ GeV and $\sqrt{s}=2.3$ GeV in both cross sections.
@@
Figure 8: (Color online) The same as Fig. 5 but in Model B.
Table 3: The fitted parameters of the process $K^{-}p\rightarrow K\Xi(1530)$, where increasing the weight (eight times) of six experimental data with high cross sections, and delete the one experimental data with low cross section in the range of $\sqrt{s}=[2.087,2.168]$ GeV.
Coupling constant Value Coupling constant Value
$g^{\prime}_{\Lambda(1116)}$ $1.116^{+0.220}_{-0.220}$ $g^{\prime}_{\Sigma(1193)}$ $-3.321^{+0.145}_{-0.145}$
$g^{\prime}_{\Lambda(1405)}$ $-0.192^{+0.085}_{-0.085}$ $g^{\prime}_{\Sigma(1670)}$ $-2.065^{+0.281}_{-0.281}$
$g^{\prime}_{\Lambda(1890)}$ $0.187^{+0.600}_{-0.600}$ $g^{\prime}_{\Sigma(2080)}$ $-0.805^{+0.217}_{-0.217}$
$g^{\prime}_{\Lambda(2110)}$ $0.545^{+0.015}_{-0.015}$ $g^{\prime}_{\Sigma(2250)}$ $0.212^{+0.034}_{-0.034}$
$g^{\prime}_{\Lambda(2325)}$ $-0.197^{+0.015}_{-0.015}$
@@
In addition to the total cross sections, the individual contributions of each intermediate state for $K^{-}p\rightarrow K^{+}\Xi(1530)^{-}$ are presented in Fig. 4. From Fig. 4-(a) one can find that the cross sections resulted from $\Sigma(1193)$ intermediate process are predominant, while the cross section resulted from $\Lambda(1116)$ is several times smaller than that of $\Sigma(1193)$. As indicated by Fig. 4-(b), the cross sections resulted from $\Lambda(1405)$ are at least 3 orders smaller than those from $\Lambda(1116)$, which indicate that the contributions from $\Lambda(1405)$ are negligible even if one considers the interference contributions with other states. In Fig. 4-(c), we present the contributions from $\Lambda(1890)$, $\Lambda(2325)$, $\Sigma(1670)$, and $\Sigma(2080)$, whose total angular momenta are $3/2$. From this diagram, one can find that the contribution from $\Lambda(1890)$ is not obvious with its contribution to the cross sections to be less than 2 $\mu$b. Meanwhile, the peak values of cross sections for $\Lambda(2325)$, $\Sigma(1670)$, and $\Sigma(2080)$ are about $20\sim$ $\mu$b. In Fig. 4-(d), we present the individual contributions from two high-spin intermediate state, i.e. $\Lambda(2110)$ and $\Sigma(2250)$. The maximum of the cross sections resulted from $\Lambda(2110)$ is estimated to be around 50 $\mu$b at $\sqrt{s}=2.13$ GeV, and the contribution from $\Sigma(2250)$ is about 1/3 of that from $\Lambda(2110)$ in the peak position.
In addition, the individual contributions from hyperon for $K^{-}p\rightarrow K^{0}\Xi(1530)^{0}$ are presented in Fig. 5. Comparing with Fig. 4, one can find that the cross sections resulted from four $\Sigma$ intermediate states are basically consistent in these two processes, which is consistent with isospin symmetry expectation. As for $\Lambda$ intermediate process, the $u$-channel contributions vanish for the $K^{-}p\to K^{0}\Xi(1530)^{0}$ process due to charge conservation. Comparing Fig. 4 and Fig. 5, one can find that the contributions from $\Lambda(1116)$ $\Lambda(1405)$ and $\Lambda(1890)$ are much larger in $K^{-}p\to K^{+}\Xi(1530)^{-}$ process, which indicates that the $u$-channel is dominant for the below threshold intermediate states. Moreover, the contributions from the above threshold $\Lambda$ resonances, such as $\Lambda(2110)$ and $\Lambda(2325)$ are similar, indicating that the $s$-channel is dominant.
@@ III.1.2 Model B
As we clarified at the beginning of this subsection, the experimental data for the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ in the range of $\sqrt{s}=[2.087,2.168]$ GeV are inconsistent with each other. In Model B, we first remove the data at $\sqrt{s}=2.11$ GeV and perform a weighted fit with the weights of the other data in the range enhanced. For example, when we set the weight of these data to 8, we can obtain the line shape of the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and reproduce the data in the range of $\sqrt{s}=[2.087,2.168]$ GeV. In this case the fitted parameters are collected in Table 2. With these parameters, the $\chi^{2}$ and weight $\chi^{2}$ is estimated to be 96.24 and 138.93, which are much larger than those in model A.
@@
Using the parameters in Table 2, we can obtain the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$, which are presented in Fig .6. Notably, the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ approach 70 $\mathrm{\mu b}$ in the vicinity of 2.1 GeV, which are consistent with the lower limit of the experimental data. Nonetheless, the fitted cross sections overestimate the experimental measurements outside the range of $\sqrt{s}=[2.087,2.168]$ GeV. For the $K^{-}p\to K^{0}\Xi(1530)^{0}$ process, the fitted curve exhibits a modest overestimation of the experimental data in the vicinity of 2.1 GeV, followed by an underestimation around 2.3 GeV. Based on a comparative analysis of Figs. 4 and 6, we conclude that Model A demonstrates a better global agreement with the experimental data than Model B.
@@
Figure 9: (Color online) The differential cross sections for $K^{-}p\rightarrow K^{+}\Xi(1530)^{-}$ depending on $\cos\theta$ with several typical values of $\sqrt{s}$. The black solid curves are obtained with the fitted parameters in Table 2, which correspond to Model A. The black dotted curves are obtained with the fitted parameters in Table 3, which correspond to Model B.
Figure 10: (Color online) The same as Fig. 9 but for the process $K^{-}p\rightarrow K^{0}\Xi(1530)^{0}$.
@@
In Figs. 7 and 8, we present the individual contributions from different intermediate states for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$, respectively. From the individual contributions, one can find that that the resonances $\Sigma(2080)$, $\Sigma(1670)$ and $\Lambda(2110)$ significantly contribute to the peak structure around 2.1 GeV. In order to enhance the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ in the vicinity of 2.1 GeV, the contributions from these resonances should be increased. By comparing the fitted parameters in Tables 2 and 3, one can find that in Model B the absolute value of the parameters $g_{\Sigma(1670)}$ and $g_{\Lambda(2110)}$ increase by $33.6\%$ and $20.0\%$, respectively, while the absolute value of $g_{\Sigma(2080)}$ decreases by $6.8\%$. However, as we analyzed in Model A, the contributions for $\Sigma$ resonances are the same for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$. In addition, the contributions from above threshold $\Lambda$ resonance are $s$-channel dominant, indicating the above threshold $\Lambda$ resonances have similar contributions to $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$ processes. Thus, the overall enhancement of the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ in the vicinity of 2.1 GeV inevitably also enhance the cross sections for $K^{-}p\to K^{0}\Xi(1530)^{0}$, which lead to a worse global agreement with the experimental data of the cross sections for $K^{-}p\to K^{0}\Xi(1530)^{0}$ near 2.1 GeV.

### v2

@@ III NUMERICAL RESULTS AND DISCUSSIONS
It is worth mentioning that a lot of $\Lambda$ and $\Sigma$ resonances have been observed below the threshold of $K\Xi(1530)$, and these states dominantly contribute to the $u$ channel of the reaction process, and the lineshape of the cross section resulted from these below-threshold states are very similar, thus, it is very difficult to determine the individual contributions from these below-threshold states by fitting the cross sections for $K^{-}p\to K\Xi(1530)$. In the present fit, we only consider $\Lambda(1116)$, $\Lambda(1405)$, $\Lambda(1890)$, $\Sigma(1193)$ and $\Sigma(1670)$, whose $J^{P}$ quantum numbers are $1/2^{\pm}$ or $3/2^{\pm}$. As for the above-threshold states, the contributions from the $s$ channel are dominant, by comparing the structure of the cross sections Berge:1966zz ; Dauber:1969hg ; Briefel:1977bp ; Flaminio:1979iz , we take $\Sigma(2080)$, $\Lambda(2110)$, $\Sigma(2250)$, and $\Lambda(2325)$ into consideration. The resonance parameters involved in the present fit are collected in Table 1. As indicated in Ref. Sharov:2011xq , for the above-threshold resonance with high-spin, the $s$-channels are dominant compared to the $u$-channel, thus, in the present estimations, only $s$-channels are considered for $\Lambda(2110)$ and $\Sigma(2250)$. Therefore, the amplitude for $K^{-}p\to K^{+}\Xi(1530)^{-}$ reads,
@@
With the above amplitudes, one can obtain the differential cross sections for the $K^{-}p\to K\Xi(1530)$ process,
@@
where $s$ refers to the square of the center-of-mass energy, and $\theta$ is the scattering angle, which is the angle of outgoing $\Xi(1530)$ and the incident kaon beam direction in the center-of-mass frame. The $\vec{q_{f}}$ and $\vec{q_{i}}$ stand for the three-momentum of the final $\Xi(1530)$ and the initial kaon beam in the center-of-mass system, respectively, and their values are presented in Eqs. (25).
@@
where $Y$ refers to the $\Lambda$ and $\Sigma$ hyperons and their resonances. In the present estimations, the values of $g_{Y}$ will be determined by simultaneously fitting the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$ .
Figure 3: (Color online) The cross sections for $(a)$ $K^{-}p\rightarrow K^{+}\Xi(1530)^{-}$ and $(b)$ $K^{-}p\rightarrow K^{0}\Xi(1530)^{0}$ depending on the $\sqrt{s}$. The black points with error bars correspond to the experimental data of the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$ from Refs. Berge:1966zz ; Dauber:1969hg ; Briefel:1977bp ; Flaminio:1979iz , while the blue solid curves are obtained with the fitted parameters in Table 2.
@@
Figure 5: (Color online) The individual contributions from different intermediate states for the cross sections for $K^{-}p\rightarrow K^{0}\Xi(1530)^{0}$ depending on the $\sqrt{s}$ in model A. It should be noted that the diagrams $(c)$ and $(d)$ are almost the same as those in Fig. 4 since the above-threshold $\Lambda$ resonance contributions are dominated by the s channel, while the $\Sigma$ resonance contributions are exactly the same in both processes.
@@ III.1 Cross sections for $K^{-}p\to K\Xi(1530)$
In the present fitting process, nine $\Lambda$ and $\Sigma$ hyperons and their resonances are taken into consideration. In addition, two kinds of cutoff constants, $\Lambda_{a}$ for low-spin intermediate states ($J\leq 3/2$) and $\Lambda_{b}$ for high-spin intermediate states ($J\geq 5/2$), are involved by the form factors in the amplitudes. Here, we take $\Lambda_{a}=1.0$ GeV and $\Lambda_{b}=0.5$ GeV as those in Ref. Sharov:2011xq , while the remaining parameters $g_{Y}$ can be determined by fitting the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$ simultaneously. In addition, it should be noted that the experimental data for the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ in the range of $\sqrt{s}=[2.087,2.168]$ GeV are inconsistent with each other as shown in Fig. 3. In particular, the cross sections at $2.087$, $2.111$, $2.123$, $2.126$, $2.151$, and $2.168$ GeV are around 80 $\mathrm{\mathrm{\mu}b}$ with large uncertainties, while the cross section at $2.108$ GeV is measured to be $(41\pm 10)\ \mathrm{\mu b}$, which is evidently smaller than the nearby data. Due to the inconsistency mentioned above, we employ two distinct fitting strategies, a uniform weighting scheme (model A) and different weighting approach (model B).
Table 2: The fitted parameters of the process $K^{-}p\rightarrow K\Xi(1530)$, where the weight of each experimental value remains the same.
Coupling Value Coupling Value
$g_{\Lambda(1116)}$ $1.37\pm+0.20$ $g_{\Sigma(1193)}$ $3.18\pm 0.07$
$g_{\Lambda(1405)}$ $0.05\pm 0.08$ $g_{\Sigma(1670)}$ $1.55\pm 0.20$
$g_{\Lambda(1890)}$ $0.50\pm 0.27$ $g_{\Sigma(2080)}$ $0.86\pm 0.20$
$g_{\Lambda(2110)}$ $0.45\pm 0.02$ $g_{\Sigma(2250)}$ $0.21\pm 0.03$
$g_{\Lambda(2325)}$ $0.18\pm 0.02$
@@
Figure 6: (Color online) The same as Fig. 3 but in model B. The points with error bars correspond to the experimental data of the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$ from Refs. Berge:1966zz ; Dauber:1969hg ; Briefel:1977bp ; Flaminio:1979iz , while the blue solid curves are obtained with the fitted parameters in Table 3. The green data are excluded in the present fit, while the weight of the six red data points in the range $\sqrt{s}=[2.087,2.168]$ GeV are set to 8.
@@ III.1.1 Model A
[…]
@@
Figure 8: (Color online) The same as Fig. 5 but in model B. The diagrams $(c)$ and $(d)$ are almost the same as those in Fig. 7 for the same reason clarified in Fig. 5.
Table 3: The fitted parameters of the process $K^{-}p\rightarrow K\Xi(1530)$, where the weight of six experimental data values with high cross sections was increased (eight times), and the one experimental data value with low cross section in the range $\sqrt{s}=[2.087,2.168]$ GeV was deleted.
Coupling Value Coupling Value
$g^{\prime}_{\Lambda(1116)}$ $1.12\pm+0.22$ $g^{\prime}_{\Sigma(1193)}$ $3.32\pm 0.15$
$g^{\prime}_{\Lambda(1405)}$ $0.19\pm 0.09$ $g^{\prime}_{\Sigma(1670)}$ $2.07\pm 0.28$
$g^{\prime}_{\Lambda(1890)}$ $0.19\pm 0.60$ $g^{\prime}_{\Sigma(2080)}$ $0.81\pm 0.22$
$g^{\prime}_{\Lambda(2110)}$ $0.55\pm 0.02$ $g^{\prime}_{\Sigma(2250)}$ $0.21\pm 0.03$
$g^{\prime}_{\Lambda(2325)}$ $0.20\pm 0.02$
@@
[…]
In addition, the individual contributions from hyperons for $K^{-}p\rightarrow K^{0}\Xi(1530)^{0}$ are presented in Fig. 5. Comparing with Fig. 4, one can find that the cross sections resulted from four $\Sigma$ intermediate states are basically consistent in these two processes, which are consistent with isospin symmetry expectation. As for the $\Lambda$ intermediate process, the $u$-channel contributions vanish for the $K^{-}p\to K^{0}\Xi(1530)^{0}$ process due to charge conservation. Comparing Fig. 4 and Fig. 5, one can find that the contributions from $\Lambda(1116)$, $\Lambda(1405)$, and $\Lambda(1890)$ are much larger in the $K^{-}p\to K^{+}\Xi(1530)^{-}$ process, which indicates that the $u$ channel is dominant for the below-threshold intermediate states. Moreover, the contributions from the above-threshold $\Lambda$ resonances, such as $\Lambda(2110)$ and $\Lambda(2325)$ are similar, indicating that the $s$ channel is dominant.
@@ III.1.2 Model B
As we clarified at the beginning of this section, the experimental data for the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ in the range $\sqrt{s}=[2.087,2.168]$ GeV are inconsistent with each other. In model B, we first remove the data at $\sqrt{s}=2.11$ GeV and perform a weighted fit with the weights of the other data in this range enhanced. We increase the weight of these data, and find that when we set the weight of these data to be 8, the lineshape of the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ can roughly reproduce the data in the range $\sqrt{s}=[2.087,2.168]$ GeV. In this case the fitted parameters are collected in Table 3. With these parameters, the $\chi^{2}$ and weighted $\chi^{2}$ are estimated to be 96.24 and 138.93, respectively, which are much larger than those in model A.
@@
Using the parameters in Table 3, we can obtain the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$, which are presented in Fig. 6. Notably, the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ approach 70 $\mathrm{\mu b}$ in the vicinity of 2.1 GeV, which is consistent with the lower limit of the experimental data. Nonetheless, the fitted cross sections overestimate the experimental measurements outside the range $\sqrt{s}=[2.087,2.168]$ GeV. For the $K^{-}p\to K^{0}\Xi(1530)^{0}$ process, the fitted curve exhibits a modest overestimation of the experimental data in the vicinity of 2.1 GeV, followed by an underestimation around 2.3 GeV. Based on a comparative analysis of Figs. 4 and 6, we conclude that model A demonstrates a better global agreement with the experimental data than model B. More accurate measurements of the cross sections in this energy range can be helpful for our understanding of $\Xi(1530)$ production in the $K^{-}p$ reaction , which should be accessible for J-PARC.
@@
Figure 9: The differential cross sections for $K^{-}p\rightarrow K^{+}\Xi(1530)^{-}$ depending on $\cos\theta$ with several typical values of $\sqrt{s}$. The black solid curves are obtained with the fitted parameters in Table 2, which corresponds to model A. The black dotted curves are obtained with the fitted parameters in Table 3, which corresponds to model B.
Figure 10: The same as Fig. 9 but for the process $K^{-}p\rightarrow K^{0}\Xi(1530)^{0}$.
@@
In Figs. 7 and 8, we present the individual contributions from different intermediate states for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$, respectively. From the individual contributions, one can find that the resonances $\Sigma(2080)$, $\Sigma(1670)$ and $\Lambda(2110)$ significantly contribute to the peak structure around 2.1 GeV. In order to enhance the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ in the vicinity of 2.1 GeV, the contributions from these resonances should be increased. By comparing the fitted parameters in Tables 2 and 3, one can find that in model B the absolute value of the parameters $g_{\Sigma(1670)}$ and $g_{\Lambda(2110)}$ increase by $33.6\%$ and $20.0\%$, respectively, while the absolute value of $g_{\Sigma(2080)}$ decreases by $6.8\%$. However, as we analyzed in model A, the contributions for $\Sigma$ resonances are the same for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$. In addition, the contributions from above-threshold $\Lambda$ resonance are $s$-channel dominant, indicating the above-threshold $\Lambda$ resonances have similar contributions to $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$ processes. Thus, the overall enhancement of the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ in the vicinity of 2.1 GeV inevitably also enhances the cross sections for $K^{-}p\to K^{0}\Xi(1530)^{0}$, which lead to a worse global agreement with the experimental data of the cross sections for $K^{-}p\to K^{0}\Xi(1530)^{0}$ near 2.1 GeV.